OpenStudy (anonymous):

The circuit parameters have the following values: VS=4V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V

OpenStudy (anonymous):

p1: Vo=Vs since vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)*vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) p1: 4s p2: 4-(1*1)(e^((2-1)/1)-1) === 2.281718 p3: ln(4/2*1*1+1)*1+1 === 2.0986 p4: sqrt(4/1)+1 === 3 p5: -e^((2-1)/1) === -e === -2.7182 p6:-(2-1) === -1